OpenStudy (chycora):

There are 52 white keys and 36 black keys on a standard piano. What is the probability that if a person did not look and hit the keyboard twice that he or she would first hit a white key and then hit a black key?

OpenStudy (chycora):

A.\[\frac{ 52 }{ 88 }\] B.\[\frac{ 1 }{ 1872 }\] C.\[\frac{ 169 }{ 484 }\] D.\[\frac{ 117 }{ 484 }\]

OpenStudy (chycora):

@TheWeeknd16 @thomaster @TheAsker2002 @smart8thgrader @Xx_H20_Avenger_xX @jabez177 @Here_to_Help15 @love_to_love_you

OpenStudy (chycora):

@Haseeb96

OpenStudy (anonymous):

Wrong subject

OpenStudy (chycora):

o yea srry but d is correct right

OpenStudy (anonymous):

Yeah